To find the distance to the nearest hundredth between the points (-6-5) and (-1,3) do I start by using -6 and -1 as my X and -5 and 3 as my y?
thank you

Found 2 solutions by stanbon, venugopalramana:
Answer by stanbon(48555)

Yes.
Cheers,
Stan H.

Answer by venugopalramana(3286)

To find the distance to the nearest hundredth between the points (-6-5) and (-1,3) do I start by using -6 and -1 as my X and -5 and 3 as my y?
thank you
CORRECT...
SQRT[(-6+1)^2+(-5-3)^2]=SQRT(25+64)=SQRT(89)=9.433981=9.43